Solve y^1=(2x+2y+3)/(2x+2y-1)

1
Expert's answer
2022-02-27T16:49:26-0500

"y'=\frac{2x+2y+3}{2x+2y-1}\\
y'=\frac{2x+2y-1+4}{2x+2y-1}\\
y'=1+\frac{4}{2x+2y-1}\\
2x+2y-1=z(x)\\
y=\frac{1}{2}(z-2x+1)\\
y'=\frac{1}{2}(z'-2)\\
\frac{1}{2}(z'-2)=1+\frac{4}{z}\\
\frac{1}{2}z'=\frac{2z+4}{z}\\
\frac{zdz}{z+2}=4dx\\
\int{\frac{zdz}{z+2}}=\int4dx\\
\frac{z+2-2dz}{z+2}=\int4dx\\
\int(1-\frac{2}{z+2})dz=4x+C\\
z-2ln|z+2|=4x+C\\
2x+2y-1-2ln|2x+2y+1|=4x+C\\
-2x+2y-1-2ln|2x+2y+1|=C"